A psychologist studied self-esteem scores and found the data set to be normally distributed with a mean of 50 and a standard deviation of 6.
1: What is the z-score associated with a raw score of 47?
2: A raw score of 47 is associated with what percentile?
3: What are the z-scores that mark the middle 95% of this distribution?
4: What is the raw score below the mean?
5: What is the raw score above the mean?
6: What is the median of this distribution?

Answer by stanbon(75887)

A psychologist studied self-esteem scores and found the data set to be normally distributed with a mean of 50 and a standard deviation of 6.
1: What is the z-score associated with a raw score of 47?
z(47) = (47-50)/6 = -3/6 = -1/2
----------------------------------------------
2: A raw score of 47 is associated with what percentile?
Ans:: P(x < 47) = P(z < -0.5) = normalcdf(-100,-0.5) = 0.3085 = 30.85%
--------------------------------------------------------------------
3: What are the z-scores that mark the middle 95% of this distribution?
Find the z-values which have a right or a left tail of 2.5%
z = -1.96 and z = 1.96
---------------------------------------
4: What is the raw score below the mean?
x = -1.96*6 + 50 = 38.24
---------------------------------------
5: What is the raw score above the mean?
x = +1.96*6 + 50 = 61.76
----------------------------------------
6: What is the median of this distribution?
median = 50
------------------
Cheers,
Stan H.
-------------